# Attention Is Not What You Need: Revisiting Multi-Instance Learning for Whole Slide Image Classification

### v1

@@ Abstract
Although attention-based multi-instance learning algorithms have achieved impressive performances on slide-level whole slide image (WSI) classification tasks, they are prone to mistakenly focus on irrelevant patterns such as staining conditions and tissue morphology, leading to incorrect patch-level predictions and unreliable interpretability. Moreover, these attention-based MIL algorithms tend to focus on salient instances and struggle to recognize hard-to-classify instances. In this paper, we first demonstrate that attention-based WSI classification methods do not adhere to the standard MIL assumptions. From the standard MIL assumptions, we propose a surprisingly simple yet effective instance-based MIL method for WSI classification (FocusMIL) based on max-pooling and forward amortized variational inference. We argue that synergizing the standard MIL assumption with variational inference encourages the model to focus on tumour morphology instead of spurious correlations. Our experimental evaluations show that FocusMIL significantly outperforms the baselines in patch-level classification tasks on the Camelyon16 and TCGA-NSCLC benchmarks. Visualization results show that our method also achieves better classification boundaries for identifying hard instances and mitigates the effect of spurious correlations between bags and labels.
@@ 1 Introduction
Whole slide images (WSIs) analysis has been widely utilized in clinical applications such as computer-aided tumour diagnosis and prognosis (Lu et al. 2021; Yao et al. 2020; Chen et al. 2021; Li, Li, and Eliceiri 2021). As a single WSI often contains billions of pixels and obtaining patch-level annotations is labour-intensive (Lu et al. 2021), it is a standard practice to crop the WSI into smaller patches and utilize Multi-Instance Learning (MIL) for classification. Multi-Instance Learning (MIL) is a weakly-supervised learning paradigm designed to learn from samples represented as bags (slides) of instances (patches). As computer-aided diagnosis models need to provide both slide-level and patch-level predictions of tumour tissues, MIL is particularly well-suited. (Li, Li, and Eliceiri 2021; Qu et al. 2022).
Figure 1: (a) MIL algorithms based on attention may produce incorrect patch-level attention map with correct slide-level prediction. (b) A causal graph for the generative procedure of patches and their labels.
@@
Most existing MIL-based WSI classification algorithms focus on slide-level predictions and utilize attention (Vaswani et al. 2017) for aggregating patch-level representations into slide-level (Ilse, Tomczak, and Welling 2018; Li, Li, and Eliceiri 2021; Shao et al. 2021; Qu et al. 2022; Zhang et al. 2022a, 2023; Tang et al. 2023, 2024; Keum et al. 2024). Generally speaking, these algorithms utilize the attention mechanism to assign weights for each instance within a bag and aggregate their weighted representations for classification. However, many attention-based MIL methods deviate from the standard MIL assumption, which can limit their interpretability and trustworthiness in medical diagnosis (Raff and Holt 2023).
The standard MIL assumption states that a multi-instance bag is positive if and only if at least one of its instances is positive (Foulds and Frank 2010; Dietterich, Lathrop, and Lozano-Pérez 1997), which naturally aligns with the goal of WSI classification: a slide should be labelled as cancerous if and only if it contains at least one cancerous patch. Unfortunately, attention-based MIL algorithms do not pose additional constraints to enforce the standard MIL assumption when aggregating the attention weights. As a result, these models can sometimes misinterpret non-causal or irrelevant instances as influential, leading to less accurate and less interpretable predictions in medical diagnoses.
@@
Another desideratum of WSI classification is that models should base their patch-level predictions on factors causally related to the tumour cells rather than spurious correlations. Fig. 1(b) illustrates the causal structure underlying the generation process of a patch $\bm{x}$, its ground-truth label $y$, and its predicted label $\hat{y}$. Three sets of factors determine the features of a patch: $\bm{Z}_{c}$, which causally influence whether a cell is cancerous, such as cellular morphology; $\bm{Z}_{s}$, which dictate stylistic elements, such as the positioning, size, and compression state of cells within the patch; and $\bm{Z}_{B}$, which are derived from the whole slide, including hematoxylin and eosin staining biases and tissue structures. Although an ideal model should make predictions based on $\bm{Z}_{c}$, most current models do not distinguish between $\bm{Z}_{c}$ and $\bm{Z}_{B}$.
Attention-based MIL methods do not enforce the standard multi-instance assumption or the causal relationship between patch-level and slide-level labels. First, as illustrated in Figure 1(a), higher attention scores can be assigned to regions that are irrelevant to the tumour, significantly diminishing the interpretability of assistive clinical diagnostics (Lin et al. 2023). Second, the models may struggle to identify tumour cells (hard instances) that are morphologically similar to normal cells and less malignant (Jögi et al. 2012; Zhang et al. 2023; Tang et al. 2023), leading to missed diagnoses in early-stage cancer patients. This is because attention-based models excessively prioritize salient instances (Tang et al. 2023; Zhang et al. 2023) and tend to exploit spurious correlations to fit the more challenging slides during training.
This paper studies the weakly supervised WSI classification problem without relying on the attention mechanism and proposes a simple yet effective instance-based MIL method: FocusMIL. Our approach focuses on the causal factors $\bm{Z}_{c}$ for predicting the label of instances while ignoring non-causal factors such as stylistic elements $\bm{Z}_{s}$ and bag-inherited information $\bm{Z}_{B}$. FocusMIL comprises a feed-forward neural network, reparameterization operations, and max-pooling. We emphasize that respecting the multi-instance assumption inherently constrains the model to classify based on tumour morphology $\bm{Z}_{c}$. The key distinction of our method from traditional max-pooling-based MIL classification methods is that we use mini-batch gradient decent with a batch size greater than one instead of SGD with a batch size of one. We discuss how a larger batch size enables the model to build better classification boundaries for recognizing hard instances and contributes to a more stable and efficient disentangling of $\bm{Z}_{c}$ and $\bm{Z}_{B}$.
Results on the Camelyon16 and TCGA-NSCLC datasets show that our approach significantly outperforms existing baselines at patch-level classification tasks and performs comparably with baselines on the slide level. Using slide-level supervision, our patch-level predictions predict salient and hard instances more accurately than existing methods. Furthermore, we designed a semi-synthetic dataset to test whether mainstream WSI classification algorithms respect the standard MIL assumptions and the results indicate that all the tested bag-level MIL methods failed this test. The code can be accessed from the supplementary material, which will be made publicly available after publication.
Figure 3: Overview of our FocusMIL with 2 slides in a mini-batch for illustration. Features are extracted from patches using a pre-trained feature extractor. The instance features are encoded into latent representations via a variational encoder. The classifier $P(Y|\bm{z})$ uses the inferred latent representations $\boldsymbol{z}$ to obtain patch-level prediction scores. Max-pooling is applied to the instance predictions of each MIL bag to derive the slide-level prediction.
@@ 2.1 Attention-based MIL for WSI classification
Most MIL-based WSI classification algorithms aggregate patch-level features into bag-level features according to their attention scores. An early bag-based method, ABMIL (Ilse, Tomczak, and Welling 2018), utilizes a trainable network to calculate the attention scores of instance features for bag-level classification and obtains a bag-level representation through their weighted sum. Subsequent methods have inherited similar ideas and extensively incorporated various attention mechanisms (Li, Li, and Eliceiri 2021; Lu et al. 2021; Shao et al. 2021; Zhang et al. 2022a).
@@
Attention-based MIL methods face the challenge of overfitting from three sources (Zhang et al. 2022a; Lin et al. 2023; Tang et al. 2023; Zhang et al. 2023). Firstly, many cancerous slides contain salient instances, i.e., areas with highly differentiated cancer cells that significantly differ from the normal ones (Qu et al. 2024; Jögi et al. 2012). Secondly, other slides may have hard instances, e.g., small cancerous areas where the cancer cells closely resemble normal cells (Bejnordi et al. 2017). Thirdly, the digital scanning equipment, slide preparation, and staining processes often introduce staining bias into WSI datasets (Zhang et al. 2022c). Existing studies (Tang et al. 2023; Zhang et al. 2023; Keum et al. 2024) have indicated that attention-based WSI classification methods may excessively focus on salient instances and fail to identify hard instances. For positive slides containing only hard instances, these methods often rely on spurious correlations caused by staining bias to fit the training set (Lin et al. 2023). Recent methods like WENO (Qu et al. 2022), MHIM (Tang et al. 2023), IBMIL (Lin et al. 2023), ACMIL (Zhang et al. 2023) and Slot-MIL (Keum et al. 2024) try to mitigate these problems; however, their generalization capability and patch-level classification performances are still not ideal.
2.2 Reproducibility in Multiple Instance Learning
@@
Recent studies (Raff and Holt 2023) have shown that many widely used MIL methods, such as MI-Net (Wang et al. 2018) and TransMIL (Shao et al. 2021), do not adhere to the standard MIL assumptions and may leverage inductive biases, such as the absence of certain instances, as signals for predicting positive bags. As staining bias is prevalent in WSI analysis (Lin et al. 2023; Zhang et al. 2022c), models that do not respect the standard MIL assumptions can learn to use these spurious correlations (Raff and Holt 2023).
@@
mi-Net (Zhou and Zhang 2002; Wang et al. 2018), and the recently proposed CausalMIL (Zhang et al. 2022b) are perhaps the only two methods that respect the MIL assumptions. CausalMIL applies a variational autoencoder (Kingma and Welling 2013) to the MIL problem, using a non-factorized prior distribution conditioned on bag information to provide a theoretical guarantee for identifying latent causal representation. However, mi-Net does not perform optimally in WSI classification, and CausalMIL has not yet been studied for WSI classification.
3 Method
3.1 Preliminaries
@@
A WSI dataset with $n$ slides are treated as MIL bags =$\mathcal{B}=\{\mathbf{B}_{1},\cdots,\mathbf{B}_{i},\cdots,\mathbf{B}_{n}\}$, where each slide has a slide-level label $Y_{i}\in\{0,1\}$ during training. Each slide is then cropped into patches corresponding to MIL instances $\mathbf{B}_{i}=\{\boldsymbol{p}_{i1},\cdots,\boldsymbol{p}_{ij},\cdots,\boldsymbol{p}_{in_{i}}\}$, where $n_{i}$ is the number of patches in the slide. For each patch $\boldsymbol{p}_{ij}$, there exists a patch-level label $y_{ij}\in\{0,1\}$ that is unknown to the learner. The standard multi-instance assumption states that a bag is positive if and only if at least one of its instances is positive:
@@
Given a feature extractor $f$, each instance is projected onto a $d$-dimensional feature vector $\boldsymbol{x}_{ij}=f(\boldsymbol{p}_{ij})\in\mathbb{R}^{d\times 1}$.
3.2 FocusMIL

### v2

# Position: From Correlation to Causation: Max-Pooling-Based Multi-Instance Learning Leads to More Robust Whole Slide Image Classification

@@ Abstract
Although attention-based multi-instance learning (MIL) algorithms have achieved impressive performance on slide-level whole slide image (WSI) classification tasks, they are prone to mistakenly focusing on irrelevant patterns such as staining conditions and tissue morphology, leading to incorrect patch-level predictions and unreliable interpretability. In this paper, we analyze why attention-based methods tend to rely on spurious correlations in their predictions. Furthermore, we revisit max-pooling-based approaches and examine the reasons behind the underperformance of existing methods. We argue that well-trained max-pooling-based MIL models can make predictions based on causal factors and avoid relying on spurious correlations. Building on these insights, we propose a simple yet effective max-pooling-based MIL method (FocusMIL) that outperforms existing mainstream attention-based methods on two datasets. In this position paper, we advocate renewed attention to max-pooling-based methods to achieve more robust and interpretable predictions.
Multi-Instance Learning, Reproducibility
@@ 1 Introduction
Figure 1: (a) A causal graph illustrating the generative procedure of patches and their labels. Ideally, $\bm{Z}_{C}$ is used to predict $y$. (b) The architecture of attention-based MIL (excluding the bag feature for brevity), where the classifier can leverage any of $\bm{Z}_{C}$, $\bm{Z}_{S}$, and/or $\bm{Z}_{B}$ for prediction.
@@
Whole slide images (WSIs) analysis has been widely utilized in clinical applications such as computer-aided tumour diagnosis and prognosis (Lu et al., 2021; Yao et al., 2020; Chen et al., 2021; Li et al., 2021). As a single WSI often contains billions of pixels and obtaining patch-level annotations is labour-intensive (Lu et al., 2021), it is a standard practice to crop the WSI into smaller patches and utilize Multi-Instance Learning (MIL) for classification. Multi-Instance Learning is a weakly-supervised learning paradigm designed to learn from samples represented as bags (slides) of instances (patches). As computer-aided diagnosis models must offer interpretable predictions at both the slide and patch levels to support clinical decision-making, MIL is particularly well-suited (Li et al., 2021; Qu et al., 2022).
@@
Most existing MIL-based WSI classification algorithms focus on slide-level predictions, using attention (Vaswani et al., 2017) for aggregating patch-level representations (Ilse et al., 2018; Li et al., 2021; Shao et al., 2021; Qu et al., 2022; Zhang et al., 2022a, 2023; Tang et al., 2023, 2024; Keum et al., 2024). As Figure 1 shows, they assign attention weights to instances within a bag and aggregate their weighted instance representations to a bag representation for classification. However, recent studies (Raff & Holt, 2023) have shown that most attention-based MIL methods deviate from the standard MIL assumption, and are not learnable at the instance-level (Jang & Kwon, ).
The standard MIL assumption states that a multi-instance bag is positive if and only if at least one of its instances is positive (Foulds & Frank, 2010; Dietterich et al., 1997). This assumption aligns well with the nature of cancer in WSI, where a slide is considered malignant if cancerous tissue or cells are present, even in a small region. Unfortunately, attention-based bag-level MIL methods can gather information from all the instances in the bag to make holistic discrimination, thereby diverging from this assumption. The violations of this assumption by attention-based methods may limit their interpretability and trustworthiness in medical diagnosis (Raff & Holt, 2023).
@@
Figure 1(a) illustrates a plausible causal structure underlying the generation process of a patch $\bm{x}$, its ground-truth label $y$, and its predicted label $\hat{y}$. We posit that patch features in WSIs are determined by three sets of factors: $\bm{Z}_{c}$, which causally influence whether a cell is cancerous, such as cellular morphology; $\bm{Z}_{s}$, which dictate stylistic elements, such as the positioning, size, and compression state of cells within the patch; and $\bm{Z}_{B}$, which are derived from the whole slide, including tissue structures, hematoxylin and eosin staining conditions, and additional biases or artifacts introduced during slide preparation. Although an ideal model should make predictions based on $\bm{Z}_{c}$, most current models do not distinguish between $\bm{Z}_{c}$ and $\bm{Z}_{B}$.
Attention-based models form bag-level features by weighting instances within each bag, as shown in Figure 1(b). This process can inadvertently leverage the bias-related factor $\bm{Z}_{B}$ to fit the training set by aggregating the instances within a bag with weighted attention rather than focusing on the true causal indicator $\bm{Z}_{C}$. As a result, these models can misinterpret non-causal or irrelevant instances as influential and produce incorrect predictions when the biases in unseen slides differ from those in the training set, undermining both robustness and interoperability. Moreover, attention-based models often miss tumour cells that are morphologically similar to normal cells and less malignant (Jögi et al., 2012; Zhang et al., 2023; Tang et al., 2023), leading to potential missed diagnoses in early-stage cancer. Their tendency to prioritize salient instances (Tang et al., 2023; Zhang et al., 2023) and rely on spurious correlations to fit the more challenging slides during training exacerbates these issues.
Although max-pooling-based methods inherently adhere to the standard MIL assumption (Raff & Holt, 2023) and have been theoretically shown to be learnable at the instance level (Jang & Kwon, ), they have been largely overlooked in recent years due to their relatively weaker classification performance. This naturally raises the question: why do max-pooling methods, which appear well-suited for WSI classification (where the presence of a single cancerous instance determines the bag label), perform poorly? Moreover, can they be refined to leverage their inherent advantages for more robust and interpretable WSI classification?
This position paper starts by empirically demonstrating that max-pooling-based methods can sometimes fail to learn discriminative features when the positive concept is difficult to capture, often overfitting the training set by memorizing the instances within the bag and performing poorly on the test set. This issue is particularly pronounced when max-pooling is combined with a feed-forward neural network. Motivated by these insights, we propose FocusMIL, a simple yet effective max-pooling-based approach that encodes latent variables into a multivariate Gaussian distribution to mitigate naive memorization. FocusMIL integrates a feed-forward neural network, reparameterization operations, and a max-pooling operator. The key differences of our method from traditional max-pooling-based MIL classification methods are: 1) learning a distribution over latent representations rather than a deterministic one, and 2) using mini-batch gradient descent with a batch size greater than one instead of a batch size of one. We discuss how the latent distribution and a larger batch enable the model to better recognize hard instances and disentangle $\bm{Z}_{c}$ and $\bm{Z}_{B}$.
We validate our position with comprehensive evaluations of attention-based and max-pooling-based algorithms on the Camelyon16 (Bejnordi et al., 2017a) and Camelyon17 (Bejnordi et al., 2017b) datasets. The results show that the max-pooling-based FocusMIL achieves comparable or better performance than state-of-the-art attention-based MIL methods at the slide-level, while significantly outperforming attention-based methods in patch-level predictions and tumour region localization. Furthermore, our experimental results on the semi-synthetic dataset confirm that the performance of attention-based methods degrades significantly when the bag context distribution differs substantially between the training and test sets, whereas max-pooling-based methods can still achieve strong results. Therefore, we argue for a renewed focus on max-pooling-based methods, emphasizing that they can offer not only improved predictive performance but also more robust and interpretable outputs for real-world clinical applications. The key insights of this paper are summarized as follows:
  1. 1.
    We discuss why attention-based methods are prone to relying on spurious correlations, undermining the reliability of their prediction.
  2. 2.
    We investigate traditional max-pooling-based methods and empirically analyze why they often fail to deliver competitive performance in WSI classification.
  3. 3.
    We propose FocusMIL, a simple max-pooling-based method that outperforms complex attention-based MIL methods, highlighting the need to revisit max-pooling-based approaches.
  4. 4.
    We conduct comprehensive experiments to demonstrate that max-pooling-based methods can learn causal factors, resulting in improved robustness for WSI classification tasks. The code will be made publicly available after acceptance.
@@ 2.1 Attention-based MIL for WSI classification
Most MIL-based WSI classification algorithms aggregate patch-level features into bag-level features according to their attention scores. An early bag-based method, ABMIL (Ilse et al., 2018), utilizes a trainable network to calculate the attention scores of instance features for bag-level classification and obtains a bag-level representation through their weighted sum. Subsequent methods have inherited similar ideas and extensively incorporated various attention mechanisms (Li et al., 2021; Lu et al., 2021; Shao et al., 2021; Zhang et al., 2022a).
@@
Attention-based MIL methods face the challenge of overfitting from three sources (Zhang et al., 2022a; Lin et al., 2023; Tang et al., 2023; Zhang et al., 2023). Firstly, many cancerous slides contain salient instances, i.e., areas with highly differentiated cancer cells that significantly differ from the normal ones (Qu et al., 2024; Jögi et al., 2012). Secondly, other slides may have hard instances, e.g., small cancerous areas where the cancer cells closely resemble normal cells (Bejnordi et al., 2017a). Thirdly, the digital scanning equipment, slide preparation, and staining processes often introduce staining bias into WSI datasets (Zhang et al., 2022c). Existing studies (Tang et al., 2023; Zhang et al., 2023; Keum et al., 2024) have indicated that attention-based WSI classification methods may excessively focus on salient instances and fail to identify hard instances. For positive slides containing only hard instances, these methods often rely on spurious correlations caused by staining bias to fit the training set (Lin et al., 2023). Recent methods like WENO (Qu et al., 2022), MHIM (Tang et al., 2023), IBMIL (Lin et al., 2023), ACMIL (Zhang et al., 2023) and Slot-MIL (Keum et al., 2024) try to mitigate these problems; however, their generalization capability and patch-level classification performances are still not ideal.
2.2 Reproducibility and Learnability in Multiple Instance Learning
@@
Recent studies (Raff & Holt, 2023) have shown that many widely used MIL methods, such as MI-Net (Wang et al., 2018) and TransMIL (Shao et al., 2021), do not adhere to the standard MIL assumptions and may leverage inductive biases, such as the absence of certain instances, as signals for predicting positive bags. As staining bias is prevalent in WSI analysis (Lin et al., 2023; Zhang et al., 2022c), models that do not respect the standard MIL assumptions can learn to use these spurious correlations (Raff & Holt, 2023). Kaczmarzyk et al. (2024) found that attention-based methods heavily rely on non-tumour regions for metastasis detection and often neglect small tumour regions.
@@
mi-Net (Zhou & Zhang, 2002; Wang et al., 2018), and the recently proposed CausalMIL (Zhang et al., 2022b) are perhaps the only two methods that respect the MIL assumptions. CausalMIL applies a variational autoencoder (Kingma & Welling, 2013) to the MIL problem, using a non-factorized prior distribution conditioned on bag information to provide a theoretical guarantee for identifying latent causal representation. However, mi-Net does not perform optimally in WSI classification, and CausalMIL has not yet been studied for WSI classification.
Furthermore, the latest research on the learnability (Jang & Kwon, ) of MIL has theoretically analyzed that all attention-based methods are unlearnable at the instance-level, whereas instance-based approaches are learnable. This further highlights the need for cautious use of attention-based MIL methods and advocates for the exploration of max-pooling-based approaches.
3 Preliminaries
3.1 Multi-Instance Learning(MIL)
@@
A WSI dataset with $n$ slides are treated as MIL bags =$\mathcal{B}=\{\mathbf{B}_{1},\cdots,\mathbf{B}_{i},\cdots,\mathbf{B}_{n}\}$, where each slide has a slide-level label $Y_{i}\in\{0,1\}$ during training. Each slide is then cropped into patches corresponding to MIL instances $\mathbf{B}_{i}=\{\bm{p}_{i1},\cdots,\bm{p}_{ij},\cdots,\bm{p}_{in_{i}}\}$, where $n_{i}$ is the number of patches in the slide. For each patch $\bm{p}_{ij}$, there exists a patch-level label $y_{ij}\in\{0,1\}$ that is unknown to the learner. The standard multi-instance assumption states that a bag is positive if and only if at least one of its instances is positive:
@@
Given a feature extractor $f$, each instance is projected onto a $d$-dimensional feature vector $\bm{x}_{ij}=f(\bm{p}_{ij})\in\mathbb{R}^{d\times 1}$.
3.2 Attention-Based MIL
We first briefly review attention-based methods in preparation for our subsequent analyses. These methods aim to learn a weighted aggregation of instance features so that the final bag representation highlights the most relevant patches. A pioneering work in this direction is the ABMIL framework (Ilse et al., 2018).
ABMIL computes an attention weight $\alpha_{ij}$ for each instance $\bm{x}_{ij}$ using a small neural network:
$\alpha_{ij}=\frac{\exp\Bigl{(}\mathbf{w}^{\top}\,\tanh\bigl{(}\mathbf{V}\,\bm{x}_{ij}^{\top}\bigr{)}\Bigr{)}}{\sum_{k=1}^{n_{i}}\exp\Bigl{(}\mathbf{w}^{\top}\,\tanh\bigl{(}\mathbf{V}\,\bm{x}_{ik}^{\top}\bigr{)}\Bigr{)}},$ (3)
where $\mathbf{V}$ and $\mathbf{w}$ are trainable parameters.
The bag-level feature $\bm{z}_{i}$ is then obtained by a weighted sum of instance embeddings:
$\bm{z}_{i}=\sum_{j=1}^{n_{i}}\alpha_{ij}\,\bm{x}_{ij}.$ (4)
Finally, $\bm{z}_{i}$ is fed into a classifier to predict the bag label $\hat{Y}_{i}$. This allows the model to automatically attend to the most “informative” instances in each bag.
Because it effectively integrates instance-level representations into bag-level ones, attention mechanisms have been widely adopted in numerous WSI classification algorithms (Li et al., 2021; Lu et al., 2021; Shao et al., 2021; Zhang et al., 2022a; Qu et al., 2022).
3.3 Max-Pooling-Based MIL
Max-pooling-based MIL enforces the standard MIL assumption by selecting the most “positive” instance within a bag. A representative method, mi-Net, uses a feed-forward neural network followed by a max-pooling operation (Wang et al., 2018). Given instance-level prediction scores $s$ for a bag, the bag prediction is the maximum of the instance scores:
$\hat{Y_{i}}=\max_{j}\{s_{ij}\}.$ (5)
This simple pooling operation guarantees that a bag is classified as positive if it contains at least one positive instance and as negative only if all instances are negative, aligning with the fundamental requirement of WSI classification.
4 Deep MIL for WSI Classification
4.1 Challenges
For WSI classification tasks, slide-level and patch-level predictions should ideally be based on the causal factor $\bm{Z}_{c}$. In practice, however, several challenges may arise:
  1. 1.
    Low proportion of positive patches. In WSI datasets (e.g., Camelyon16 (Bejnordi et al., 2017a)), the percentage of positive patches can be below 10%, making it difficult for models to learn the positive concept.
  2. 2.
    Variation in tumour cell morphology. For WSI classification, the morphology of tumour cells may be different across slides due to the varying malignancy levels of cancer in patients (Zhang et al., 2023; Qu et al., 2024; Jögi et al., 2012). As shown in Figure 2, some slides contain large tumour areas with significant morphological differences from normal cells. In contrast, other slides have smaller tumour areas where the tumour cells have higher differentiation and less pronounced differences from normal cells (Qu et al., 2024), which are referred to as “hard instances” in some literature (Tang et al., 2023).
  3. 3.
    Biases introduced during slide preparation. As the slice preparation process may be different across hospitals, the bag context information $\bm{Z}_{B}$ can vary significantly and may correlate with slide labels. For real-world deployment, models should exhibit strong out-of-distribution (OOD) generalization performance.
Next, we discuss how these challenges affect model behaviour in the following two subsections.
4.2 Why Attention-Based MIL Methods are Not Stable
Attention-based MIL methods inherently violate the standard MIL assumption, as the aggregated bag representation freely utilizes information from all instances within a bag (Raff & Holt, 2023). When the morphology of the tumour cells in the slide is more significant and the area is large, these models easily learn discriminative features. However, when tumour cells (hard instances) resemble normal cells, they often struggle to learn the positive concept and instead rely on $\bm{Z}_{B}$ to fit the training set, assigning high attention scores to irrelevant patches. Our visualization results in the Appendix demonstrate that these models may assign high attention scores to “blurred regions”.
Another pitfall of attention-based models prediction is that they may treat $\bm{Z}_{B}$ information from normal tissue as a signal indicating negative slide labels, leading to missed diagnoses. In our experiments on a semi-synthetic dataset, we inject poisons into normal slides in the training set, whereas in the validation set only tumour slides would contain these harmful signals. The results show that the attention-based approaches regard these poisons as a signal of negative slides, and fail entirely on the validation set. Kaczmarzyk et al. (2024) also found that compared to predicting using all patches in a slide, predicting using only tumour patches can achieve higher sensitivity, confirming that ABMIL leverages $\bm{Z}_{B}$ information from normal patches as negative signals. According to the standard MIL assumption, negative instances should not influence the bag label (Raff & Holt, 2023).
In addition, violating the standard multi-instance learning assumptions makes slide-level predictions sensitive to the number of tumour patches. Kaczmarzyk et al. (2024) demonstrated that even when using features extracted by foundation models like UNI (Chen et al., 2024), REMEDIS (Azizi et al., 2023), and CTransPath (Wang et al., 2022), the slide-level predictions are still affected by the size of the tumour regions. This could potentially lead to the disastrous consequence of missed diagnoses.
4.3 Revisiting Max-Pooling-Based MIL